Find the wavelengths of electromagnetic waves with the following frequencies. (Assume the waves are traveling in a vacuum.) 1Will rate for answer with work!
Answer #1

Part (a)

Speed of light in vacuum = c = 3 x 108 m/s

Frequency of the electromagnetic wave = f = 2.50 x 1019 Hz

Wavelength of the electromagnetic work = \lambda

c = f\lambda

\lambda = \frac{c}{f}

\lambda = \frac{3\ast 10^{8}}{2.50\ast 10^{19}}

\lambda = 1.2 x 10-11 m

1 picometer = 10-12 meter

\lambda = 12 pm

Part (b)

Speed of light in vacuum = c = 3 x 108 m/s

Frequency of the electromagnetic wave = f = 5.50 x 109 Hz

Wavelength of the electromagnetic work = \lambda

c = f\lambda

\lambda = \frac{c}{f}

\lambda = \frac{3\ast 10^{8}}{5.50\ast 10^{9}}

\lambda = 0.0545 m

\lambda = 5.45 cm

a) Wavelength of the electromagnetic wave with a frequency of 2.50 x 1019 Hz = 12 pm

b) Wavelength of the electromagnetic wave with a frequency of 5.50 x 109 Hz = 5.45 cm